OpenStudy (anonymous):

A local True Value Hardware Store marks its goods up 38% on cost. If a snowblower cost True Value $400, the selling price would be:

OpenStudy (anonymous):

$152? i don't know...

OpenStudy (anonymous):

selling price = cost + markup = 400 + 38% of 400 = 400 + 0.38*400 = 400 +152 = 552 so the selling price will be $552

OpenStudy (mathteacher1729):

QUICK PRIMER ON PERCENTS 3.6% of a number means "multiply that number by 0.036" 67% of a number means "multiply that number by 0.67" 100% of a number means "multiply that number by 1.00" 105% of a number means "multiply that number by 1.05" So this problem is saying that the original cost of this thing was marked UP by 38%. After (AFTER) the markup, it now $400. So as an equation that means: (ORIGINAL PRICE) + 0.38 times (ORIGINAL PRICE) = 400 We can use "P" for "original price" and have P + 0.38P = 400 Which is just 1.38P = 400 Can you solve for P now? :)

OpenStudy (mathteacher1729):

OH MAN I READ IT WRONG. I read it as saying "If True Value sells the thing at 400% AFTER markup,... how much was it worth BEFORE markup?" Sorry 'bout that. Harkirat has it correct.

OpenStudy (anonymous):

its ok thanks guys!!!